OpenStudy (anonymous):

I need help with : Express the domain of the given function using interval notation click and you can see more! please help!

OpenStudy (anonymous):

\[f(x)\sqrt{2x ^{2}-8}\]

OpenStudy (anonymous):

f(x)=*

OpenStudy (anonymous):

2x^2 - 8 > 0 2x^2 = 8 x^2 = 4 x =2 , -2 take values : smaller than -2 - > -3 between -2 and 2 -> 0 larger than 2 3 2(-3)^2 - 8 = 2 * 9 - 8 = 10 positive 2*(0)^2 - 8 = -8 negative 2*(3)^2 - 8 = 2*9-8 =10 positive so x<-2 or x>2 interval notation (-infinity,-2) or (2,infinity)

OpenStudy (anonymous):

thanks so much!